A monopolist firm has the following total cost function and demand function and show your calculation.

Quantity demand price total cost

5 8 20

6 7 22

7 6 25

8 5 29

9 4 34

10 3 40

Explain what price will be charged and what output will be produced.     


1
Expert's answer
2021-08-17T17:03:18-0400




"TR =P\times Q\\MR=\frac{\Delta TR}{\Delta Q}\\MC=\frac{\Delta TC}{\Delta Q}"

Profit is maximized for monopolist when MR=MC

"Condition\space is\space satisfied\space from\space the\space table\space at\space Q=6\space and\space P=7\\

MR=MC=2"

Therefore at this price and quantity, the profit for monopolist firm is maximised.

"therefore\\

monopolist\space will\space charge\space price\space P=7\space and \space produce\space output\space Q=6"